# Pair density waves in the strong-coupling two-dimensional Holstein-Hubbard model: A variational Monte Carlo study

### v2

@@ Abstract
We thus investigate an effective $t$-$J$-$V$ model on the 2D triangular lattice with up to $21\times 21$ sites using a variational Monte Carlo (VMC) approach. Specifically, we consider a set of variational states with different patterns of broken symmetries with variable strengths of uniform (BCS) superconducting, PDW, charge-density-wave, nematic, and time-reversal symmetry breaking order parameters. We use VMC to identify the lowest variational energy among all these states. Despite known shortcomings of variational approaches to many-body systems [64], such approaches often have proven successful in identifying qualitative aspects of many ground-state phase diagrams. Parameters for the effective $t$-$J$-$V$ model [Eq. (Pair density waves in the strong-coupling two-dimensional Holstein-Hubbard model: A variational Monte Carlo study) below] we have studied are derived from the original Holstein-Hubbard model (see Table 1 above).
@@
Table 1: Parameters in the effective $t$-$J$-$V$ model [Eq. (Pair density waves in the strong-coupling two-dimensional Holstein-Hubbard model: A variational Monte Carlo study)] derived from the original Holstein-Hubbard model by fixing $t<0$, $U_{\text{e-e}}/|t|=22$, and $U_{\text{e-ph}}/|t|=18$ but varying the phonon frequency $\omega_{\text{D}}$ (see Ref. [62] for details).
@@
  2.0 0.20  0.65 $-0.056$  0.0081  0.0096
  2.5 0.20  0.66 $-0.14$  0.019  0.024
  3.0 0.21  0.67 $-0.24$  0.033  0.044
The VMC approach we adopt uses Gutzwiller projected trial wave functions, which is a powerful method to deal with strongly correlated systems [65, 66]. Specifically, we use Gutzwiller-projected mean-field-type wave functions of the form $P_{\rm{N_{e}}}P_{\rm G}|\Psi_{\rm mf}\rangle$ where $P_{\rm G}$ projects out any states with doubly occupied sites, $P_{\rm{N_{e}}}$ is the projection operator onto states with a fixed number of electrons $N_{e}$, and $|\Psi_{\rm mf}\rangle$ is constructed as the ground state of the following quadratic (mean-field-like) Hamiltonian [67],
@@
where $\mathcal{F}_{i}=(c_{i\uparrow},c_{i\downarrow})^{\mathsf{T}}$, ¯$\bar{\mathcal{F}}_{i}=(c_{i\downarrow}^{\dagger},-c_{i\uparrow}^{\dagger})^{\mathsf{T}}$, and $\boldsymbol{a}_{l}$ with $l\in\{1,2,3\}$ are primitive vectors shown in Fig. 1(a). Here $T_{i,a_{l}}$ are effective complex hopping amplitudes, $\Delta_{i,a_{l}}$ effective pair-fields, $\Lambda_{i}$ is a site energy, and $\mathbf{M}_{i}$ is an effective Zeeman field that induces various patterns of magnetic order. A pattern of phase order for the hopping amplitudes is assumed, as shown in Fig. 1(a), where the arrows indicate the direction of positive $\phi_{i,a_{l}}$ in $T_{i,a_{l}}=|T_{i,a_{l}}|e^{i\phi_{i,a_{l}}}$ with $\phi_{i,a_{l}}\in(-\pi,\pi$].
The various mean-field wave functions that we have analyzed correspond to different patterns and magnitudes of the variational parameters in Eq. (Pair density waves in the strong-coupling two-dimensional Holstein-Hubbard model: A variational Monte Carlo study). A PDW state corresponds to a spatial varying pairing field,
$\Delta_{i,a_{l}}=\left\{\begin{array}[]{ll}\tilde{\Delta}_{l}(e^{i\boldsymbol{q}\cdot\boldsymbol{r}_{i}}+e^{-i\boldsymbol{q}\cdot\boldsymbol{r}_{i}}),&\text{LO-type}\vspace{2mm}\\ \tilde{\Delta}_{l}e^{i\boldsymbol{q}\cdot\boldsymbol{r}_{i}},&\text{FF-type}\end{array}\right.$ (5)
where $\tilde{\Delta}_{l}$ represents pairing field on bonds along $\boldsymbol{a}_{l}$ direction. Whenever $\boldsymbol{q}=\boldsymbol{K}$, which is the high-symmetry momentum point respecting $C_{3}$ rotational symmetry, we can set $\tilde{\Delta}_{l}=\Delta_{s}$, $\Delta_{d}\cos(2\theta_{l})$, or $\Delta_{d+id}e^{2i\theta_{l}}$ corresponding to s-wave PDW, or d-wave nematic PDW, d+id-wave PDW, respectively ($\theta_{l}$ being the azimuth angle of $\boldsymbol{a}_{l}$). The s-wave PDW is invariant under $C_{3}$, while a d-wave PDW, in which $\theta_{l}$ can be 00, $32\pi/$ 3, or $34\pi/$ 3, is necessarily nematic [68, 69, 70]. Note that $\boldsymbol{q}$ is the center of mass momentum of the pairing in the mean-field wave function before Gutzwiller projection and it is a variational parameter. A translationally-invariant SC trial wave function corresponds to $\boldsymbol{q}=0$. We also consider CDW order through modulated $\Lambda_{i}$, $120^{\circ}$-AFM order through three-sublattice local field $\mathbf{M}_{i}$, as well as states with co-existing multiple broken symmetries [for details see Supplemental Material (SM)].
@@
For completeness, we have cursorily explored the evolution of the ground-state to larger values of the electron density $n$, beyond the PDW ordering range. Here, the typical ground state is a CDW without magnetic orders. The CDW is obtained from the projection of a mean-field state with spatial variations of on-site energy $\Lambda_{i}$, as shown in Table 2. We calculated CDW states with different momenta to obtain the state with the lowest possible energy. We find that the momentum $\boldsymbol{k}^{*}$ of the optimized CDW state is typically incommensurate, with an amplitude that decreases when increasing $n$. The density structure factor $C(\boldsymbol{k})=\frac{1}{N^{2}}\sum_{\boldsymbol{r}_{i},\boldsymbol{r}_{j}}e^{i\boldsymbol{k}\cdot(\boldsymbol{r}_{i}-\boldsymbol{r}_{j})}\langle n_{\boldsymbol{r}_{i}}n_{\boldsymbol{r}_{j}}\rangle$ is indeed peaked at the incommensurate ±$\pm\boldsymbol{k}^{*}$ as well as at the $\boldsymbol{\Gamma}$ point (which represents the uniform component of the charge density), as shown in Fig. 4(a). In other words, it is an incommensurate CDW.
@@
Acknowledgement: This work was supported in part by the NSFC under Grants No. 12347107 and No. 12404170. ZH and SAK were funded in part by the Department of Energy, Office of Basic Energy Sciences, under Contract No. DE-AC02-76SF00515 at Stanford. HY was supported in part by the New Cornerstone Science Foundation through the Xplorer Prize.

### v3

# Pair-density-wave phase of strongly interacting electrons on the triangular lattice: A variational Monte Carlo study

@@ Abstract
We thus investigate an effective $t$-$J$-$V$ model on the 2D triangular lattice with up to $21\times 21$ sites using a variational Monte Carlo (VMC) approach. Specifically, we consider a set of variational states with different patterns of broken symmetries with variable strengths of uniform (BCS) superconducting, PDW, charge-density-wave, nematic, and time-reversal symmetry breaking order parameters. We use VMC to identify the lowest variational energy among all these states. Despite known shortcomings of variational approaches to many-body systems [64], such approaches often have proven successful in identifying qualitative aspects of many ground-state phase diagrams. Parameters for the effective $t$-$J$-$V$ model [Eq. (Pair-density-wave phase of strongly interacting electrons on the triangular lattice: A variational Monte Carlo study) below] we have studied are derived from the original Holstein-Hubbard model (see Table 1 below).
@@
Table 1: Parameters in the effective $t$-$J$-$V$ model [Eq. (Pair-density-wave phase of strongly interacting electrons on the triangular lattice: A variational Monte Carlo study)] derived from the original Holstein-Hubbard model by fixing $t<0$, $U_{\text{e-e}}/|t|=22$, and $U_{\text{e-ph}}/|t|=18$ but varying the phonon frequency $\omega_{\text{D}}$ (see Ref. [62] for details).
@@
  2.0 0.20  0.65 0.056-0.056  0.0081  0.0096
  2.5 0.20  0.66 0.14-0.14  0.019  0.024
  3.0 0.21  0.67 0.24-0.24  0.033  0.044
The VMC approach we adopt uses Gutzwiller projected trial wave functions, which is a powerful method to deal with strongly correlated systems [65, 66, 67]. Specifically, we use Gutzwiller-projected mean-field-type wave functions of the form $P_{\rm{N_{e}}}P_{\rm G}|\Psi_{\rm mf}\rangle$ where $P_{\rm G}$ projects out any states with doubly occupied sites, $P_{\rm{N_{e}}}$ is the projection operator onto states with a fixed number of electrons $N_{e}$, and $|\Psi_{\rm mf}\rangle$ is constructed as the ground state of the following quadratic (mean-field-like) Hamiltonian [68, 69],
@@
where $\mathcal{F}_{i}=(c_{i\uparrow},c_{i\downarrow})^{\mathsf{T}}$, $\bar{\mathcal{F}}_{i}=(c_{i\downarrow}^{\dagger},-c_{i\uparrow}^{\dagger})^{\mathsf{T}}$, and $\boldsymbol{a}_{l}$ with $l\in\{1,2,3\}$ are primitive vectors shown in Fig. 1(a). Here $T_{i,a_{l}}$ are effective complex hopping amplitudes, $\Delta_{i,a_{l}}$ effective pair-fields, $\Lambda_{i}$ is a site energy, and $\mathbf{M}_{i}$ is an effective Zeeman field that induces various patterns of magnetic order. A pattern of phase order for the hopping amplitudes is assumed, as shown in Fig. 1(a), where the arrows indicate the direction of positive $\phi_{i,a_{l}}$ in $T_{i,a_{l}}=|T_{i,a_{l}}|e^{i\phi_{i,a_{l}}}$ with $\phi_{i,a_{l}}\in(-\pi,\pi$].
The various mean-field wave functions that we have analyzed correspond to different patterns and magnitudes of the variational parameters in Eq. (Pair-density-wave phase of strongly interacting electrons on the triangular lattice: A variational Monte Carlo study). A PDW state corresponds to a spatial varying pairing field,
$\Delta_{i,a_{l}}=\left\{\begin{array}[]{ll}\tilde{\Delta}_{l}(e^{i\boldsymbol{q}\cdot\boldsymbol{r}_{i}}+e^{-i\boldsymbol{q}\cdot\boldsymbol{r}_{i}}),&\text{LO-type}\vskip 5.69054pt\\ \tilde{\Delta}_{l}e^{i\boldsymbol{q}\cdot\boldsymbol{r}_{i}},&\text{FF-type}\end{array}\right$. (5)
where $\tilde{\Delta}_{l}$ represents pairing field on bonds along $\boldsymbol{a}_{l}$ direction. Whenever $\boldsymbol{q}=\boldsymbol{K}$, which is the high-symmetry momentum point respecting $C_{3}$ rotational symmetry, we can set $\tilde{\Delta}_{l}=\Delta_{s}$, $\Delta_{d}\cos(2\theta_{l})$, or $\Delta_{d+id}e^{2i\theta_{l}}$ corresponding to s-wave PDW, or d-wave nematic PDW, d+id-wave PDW, respectively ($\theta_{l}$ being the azimuth angle of $\boldsymbol{a}_{l}$). The s-wave PDW is invariant under $C_{3}$, while a d-wave PDW, in which $\theta_{l}$ can be 0, $2\pi/3$, or $4\pi/3$, is necessarily nematic [70, 71, 72]. Note that $\boldsymbol{q}$ is the center of mass momentum of the pairing in the mean-field wave function before Gutzwiller projection and it is a variational parameter. A translationally-invariant SC trial wave function corresponds to $\boldsymbol{q}=0$. We also consider CDW order through modulated $\Lambda_{i}$, $120^{\circ}$-AFM order through three-sublattice local field $\mathbf{M}_{i}$, as well as states with co-existing multiple broken symmetries [for details see Supplemental Material (SM)].
@@
For completeness, we have cursorily explored the evolution of the ground-state to larger values of the electron density $n$, beyond the PDW ordering range. Here, the typical ground state is a CDW without magnetic orders. The CDW is obtained from the projection of a mean-field state with spatial variations of on-site energy $\Lambda_{i}$, as shown in Table 2. We calculated CDW states with different momenta to obtain the state with the lowest possible energy. We find that the momentum $\boldsymbol{k}^{*}$ of the optimized CDW state is typically incommensurate, with an amplitude that decreases when increasing $n$. The density structure factor $C(\boldsymbol{k})=\frac{1}{N^{2}}\sum_{\boldsymbol{r}_{i},\boldsymbol{r}_{j}}e^{i\boldsymbol{k}\cdot(\boldsymbol{r}_{i}-\boldsymbol{r}_{j})}\langle n_{\boldsymbol{r}_{i}}n_{\boldsymbol{r}_{j}}\rangle$ is indeed peaked at the incommensurate $\pm\boldsymbol{k}^{*}$ as well as at the $\boldsymbol{\Gamma}$ point (which represents the uniform component of the charge density), as shown in Fig. 4(a). In other words, it is an incommensurate CDW.
@@
Acknowledgement: This work was supported in part by the NSFC under Grant No. 12347107. JW acknowledges support from the NSFC under Grant No. 12404170 and the start-up grant at HZNU. ZH and SAK were funded in part by the Department of Energy, Office of Basic Energy Sciences, under Contract No. DE-AC02-76SF00515 at Stanford. HY was supported in part by the New Cornerstone Science Foundation through the Xplorer Prize.